Please help me solve the system of equations y = 2x - 10 and x - y = 7.
Thank you for the help in advance.

Answer by glabow(165)

The simplest way to solves these two equations simultaneously is to use substitution.
y = 2x - 10
x - y = 7
The second equation is rearranged to give the value of x:
x = 7 + y [add y to both sides]
Substituting the expression 7+y for x in the first equation gives
y = 2(7+y) - 10
which is
y = 14 + 2y - 10
-y = 4 [subtract 2y from both sides]
y = -4 [multiply both sides by -1]
So, x = 7+y = 7+(-4) = 3.
Checking,
-4 = 6 - 10 [first equation]
3 -(-4) = 7 [second equation]